OpenStudy (anonymous):

Find the exact circumference? R=3 1/2ft A 14ft B 7ft C 3.5 ft D 1.75ft

OpenStudy (anonymous):

if it is a circle with radius R ==7/2 ft \[circumference~ C=2*\pi*r=2*\frac{ 22 }{ 7 } *\frac{ 7 }{2 } =?\]

OpenStudy (anonymous):

13.28?

OpenStudy (gudden):

megs... if its a circle then my circumference is coming out to be 22... i think u must check the question... and the options,